OpenStudy (anonymous):

Find the area as an integral in polar coordinates. r=4sec(theta-pi/4)

OpenStudy (anonymous):

Are there bounds for \(\theta\)?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

The reason I ask is because \(r=\sec\theta\) is the equation of a line, since \[r=\sec\theta~~\iff~~r\cos\theta=1~~\iff~~x=1\] which means you're going to need some restriction for the domain, otherwise the area is infinite.

OpenStudy (anonymous):

Changing to rectangular coordinates (so that we can see what the graph looks like more easily) you would have \[\begin{align*}r&=4\sec\left(\theta-\frac{\pi}{4}\right)\\ r\cos\left(\theta-\frac{\pi}{4}\right)&=4\\ r\bigg[\cos\theta\cos\frac{\pi}{4}+\sin\theta\sin\frac{\pi}{4}\bigg]&=4\\ \frac{1}{\sqrt2}r\bigg[\cos\theta+\sin\theta\bigg]&=4\\ r\cos\theta+r\sin\theta&=4\sqrt2\\ x+y&=4\sqrt2\end{align*}\] which is a line that looks like this:  I think the shaded region is the area you want.